The diameters of oranges from a Florida orchard have a mean of 3.2 inches with a standard deviation of 1.2 inches. A packing supplier is packaging 52 oranges in a special presentation. What is the probability that the mean diameter for these oranges is more than 2.9​ inches?

What is the probability?
Answer #1

Solution :

Given that ,

mean = = 3.2

standard deviation = = 1.2

n = 52

= 3.2

= / n =1.2/ 52 = 0.17

P( > 2.9) = 1 - P( <2.9 )

= 1 - P[( - ) / < (2.9-3.2) /0.17 ]

= 1 - P(z < -1.76)

Using z table

= 1 - 0.0392

= 0.9608